# Practical multi-fidelity machine learning: fusion of deterministic and Bayesian models

### v1

@@ 4.1.2 Comprehensive experiments of KRR-LR-GPR
Considering 10 analytical functions, we train them for the previously mentioned methods and report statistical results by repeating every experiment with 10 random initializations of the hyperparameters. Two experiments are conducted to benchmark the effectiveness and efficiency of the developed KRR-LR-GPR. All 10 functions lead to similar conclusions; therefore, we arbitrarily select the Booth function for demonstration (results of other functions are reported in Appendix D.1). Fig.7 shows the results assuming a fixed number of LF samples ($200\times d$ where $d$ is the dimension of the particular function being trained), and then consider a different number of HF samples (from $5\times d$ to $30\times d$). Fig. 8 pertains to a different experiment where we fix the number of HF samples to $20\times d$, and consider different number of LF samples (from $50\times d$ to $300\times d$). Moreover, Gaussian noise with standard deviations of $\sigma^{l}=0.3$ and $\sigma^{h}=0.3$ is considered for the datasets in both cases.
@@
Figure 8: Comparison of KRR-LR-GPR with state-of-the-art methods considering different sizes of the LF dataset (Booth function, $d=2$, Pearson cor. coef. $r=0.925$ ).
@@
The figures reinforce the argument that considering KRR for the LF and GPR for the residual creates a MF model with comparable performance to considering a GPR for all fidelities (hierarchically or not), while the computational cost decreases dramatically. Essentially, the computational complexity is reduced to $\mathcal{O}((N^{h})^{3}+(N^{l})^{2})$ instead of $\mathcal{O}((N^{h}+N^{l})^{3})$. For cases where the HF training data is scarce and low dimensional, we recommend to use a Bayesian method like GPR for that fidelity and a scalable method like KRR for the LF.
@@ 4.2.1 Illustrative example
We start with a one-dimensional illustrative example (data-scarce and low-dimensional) with 11 HF samples and 201 LF samples distributed uniformly. Aiming at explaining the limitations of the different MF models with DNNs and BNNs, we select three different LF datasets, labeled as “LF data 1”, “LF data 2” and “LF data 3”, obtained by considering different transformations of the HF data according to Eq. (50). Model performance and accuracy metrics are shown in Fig. 9 and Table 4.2.1. Table 4.2.1 also includes the different Pearson correlation coefficients ($r$) between LF and HF data. The figure includes the single fidelity BNN prediction at the top left.
@@
[h] Results comparison of DNN-LR-BNN on the 1D noisy illustrative example LF Dataset Methods NRMSE R2 Score TLL BNN 1.0925 -0.8377 -4.5359 LF Data 1 ($r$ = -0.1019) DNN-BNN 0.2102 0.9319 0.7982 DNN-LR1-BNN 1.0431 -0.6755 -7.7032 DNN-LR2-BNN 0.3196 0.8426 0.4925 LF Data 2 ($r$ = 0.9999) DNN-BNN 0.2234 0.9230 1.0007 DNN-LR1-BNN 0.0854 0.9887 1.4982 LF Data 3 ($r$ = -0.0018) DNN-BNN 1.3933 -1.9895 -4.7059 DNN-LR1-BNN 1.0607 -0.7326 -5.1139
    The subscripts 1 and 2 indicate different LF basis function orders utilized by the linear regression (LR) transfer learning model.
Fig. 9 and Table 4.2.1 illustrate that if LF and HF data are not linearly correlated (as in “LF data 1” and “LF data 3” examples), then considering linear regression with first-order terms to do transfer learning between the DNN and BNN leads to a prediction by the DNN-LR-BNN model that becomes similar to the single-fidelity BNN prediction. This directly results from the LR transfer-learning model determining a low correlation (parameter $\rho_{1}$ becomes close to zero). We believe that this is a strength of the proposed strategy because it automatically informs the analyst about the correlation between LF and HF data, suppressing the LF influence on the HF prediction when this correlation does not exist, favoring more parsimonious models (closer to the single-fidelity BNN). We also include an example of a quadratic LR transfer-learning model for the “LF data 1” example, labeled as DNN-LR2-BNN, where we show that the performance improves significantly when compared to a first-order LR model (labeled as “DNN-LR1-BNN”). Evidently, if LF and HF data become better correlated (as in “LF data 2” example), the DNN-LR-BNN model tends to have better performance.
@@
Figure 9: One-dimensional illustrative example comparing the proposed DNN-LR-BNN method (right column) with a BNN trained exclusively with HF data, and with a simple DNN-BNN (left column). Three LF datasets are employed whose Person correlation coefficients are $-0.1019$, $0.9999$, and $-0.0018$ correspondingly.
In addition, we want to highlight that being tempted to use the most flexible model, i.e. DNN-BNN, also has disadvantages as it tends to a less parsimonious model that risks overfitting (see bottom left of Fig. 9). This can be dangerous in the context of multi-fidelity because the LF often includes random and systematic errors that disrupt the HF predictions. Hence, the early literature in multi-fidelity considered first-order linear relationships between LF and HF models governed by a single hyperparameter ($\rho_{1}$). In general, however, performing model selection and establishing a simple (but not simpler) transfer-learning model between the DNN and BNN is a key consideration in practice.
@@
We also want to bring to the reader’s attention that neural network training is highly dependent on architectural and hyperparametric choices. Appendix F.2 details the dependence on hyperparameters and reinforces that the DNN-LR-BNN model is simpler to train and is less sensitive to hyperparameter choices than the more flexible DNN-BNN model. We observed in all cases that the prediction of uncertainty depends on the width of the BNN, i.e. on the number of neurons in the hidden layers. A wide network tends to the same uncertainty prediction obtained by GPR, while the prediction of the expected value (the mean) converges even for shallower BNNs.
4.2.2 Higher dimensional examples
@@
Table 5 shows that the single fidelity BNN leads to poor predictions for all examples according to all accuracy metrics. Also for all cases, the DNN-LR-BNN outperforms DNN-BNN. For the 4D function, unsurprisingly, the learned parameters $\hat{\rho}_{0}=0.86$ and $\hat{\rho}_{1}=-0.41$ are very close to the ground-truth $\rho_{0}=0.83$ and $\rho_{1}=-0.33$ showed in Eq. (55).
@@ 5 Discussion and limitations
Expressive machine learning algorithms typically have limited extrapolation ability because their large number of parameters is used to explain the training date without necessarily discovering the underlying functional form that explains the data.. DNNs often make unreliable predictions for points away from sampled points; conversely, Bayesian models such as GPRs and BNNs eventually reverse to the prior [62]. A similar issue also occurs in the DNN-BNN architecture where the HF prediction is distorted in the unsampled region as reported in Fig. 10. However, KRR-LR-GPR and DNN-LR-BNN alleviate this shortcoming to some degree because they remain smooth and close to the HF truth. The proposed transfer-learning model in Eq. (4) narrows the gap between $\mathbf{m}(\mathbf{X}^{h})^{T}\bm{\rho}$ and $\mathbf{y}^{h}$, while the BNN captures the residual utilizing a prior with zero mean that reverses to another constant after training (see purple line in Fig. 10). Therefore, the HF prediction is influenced by the parsimonious (simple) transfer-learning model, having higher chances of reasonable extrapolation.
Figure 10: Equivalence of KRR-LR-GPR and DNN-LR-BNN: The LF samples locate from zero to two while the HF samples stay in the region between zero and one
However, when we compare the DNN-LR-BNN model prediction with the KRR-LR-GPR one, we see that although the mean is similar, the uncertainty prediction is not. The DNN-LR-BNN uncertainty is smaller at the extrapolation region, demonstrating the model is overconfident. We identify two reasons for this observation. First, the inference of the BNN is approximate and does not yield the exact posterior predictive distribution. This bottleneck can be resolved for a specific problem by fine-tuning the hyperparameters of BNN using test data or cross-validation, increasing the number of neurons in the hidden layers, as previously shown. Second, there is an inherent bottleneck in evaluating uncertainty via Eq. (16) when a BNN is employed for $f^{h}(\mathbf{x})$ inference: the covariance matrix $\mathbf{K(X,X)}$ is difficult to compute and so the uncertainty present in the LF model is difficult to transport to the HF prediction.

### v2

@@ 4.1.2 Comprehensive numerical experiments of KRR-LR-GPR
The methods mentioned above are trained on 10 other analytical benchmark functions, reporting results by repeating every experiment with 5 random initializations of the training data generation. All 10 functions lead to similar conclusions; therefore, we arbitrarily select the Booth function to show in the main text (results for other functions are reported in Section D.1). Figure 7 shows the results assuming a fixed number of LF samples ($200\times d$ where $d$ is the dimension of the particular function being trained), and then considering a different number of HF samples (from $5\times d$ to $30\times d$). Figure 8 pertains to a different experiment where we fix the number of HF samples to $20\times d$, and consider different numbers of LF samples (from $50\times d$ to $300\times d$). Moreover, Gaussian noise with standard deviations of $\sigma_{l}=0.3$ and $\sigma_{h}=0.3$ is considered for the datasets in both cases.
Figure 7: Comparison of KRR-LR-GPR with state-of-the-art MF methods for different sizes of HF datasets when learning the Booth function, and considering $200\times d$ LF samples (in the case of the Booth function $d=2$, and the Pearson correlation coefficient is $r=0.925$). Different colors represent different methods. To interpret the results, note that smaller NRMSE or higher TLL indicate better performance. Note the significantly faster LF training time for KRR-LR-GPR compared to others.
@@
Figure 8: Comparison of KRR-LR-GPR with state-of-the-art methods when learning the Booth function but now considering different sizes of the LF dataset while keeping the number of HF samples fixed to $20\times d$.
@@
The figures (Figure 7, Figure 8, Figure 15, Figure 16) reinforce the argument that considering KRR for the LF and GPR for the residual creates a MF model with comparable performance to considering a GPR for all fidelity levels (hierarchically or not), while the computational cost decreases dramatically. Essentially, the computational complexity is reduced to $\mathcal{O}((N^{h})^{3}+(N^{l})^{2})$ instead of $\mathcal{O}((N^{h}+N^{l})^{3})$. For cases where the HF training data is scarce and low dimensional, we recommend using a Bayesian method like GPR for that fidelity and a scalable method like KRR for the LF. It is noted that we also conducted experiments with different HF noise levels in Section E.3 as well as an ablation study on changing both LF and HF samples in Appendix D, those results also support the above arguments.
4.1.3 Engineering application: NACA 0012 airfoil
The KRR-LR-GPR method is also used to determine the aerodynamic coefficients of a NACA 0012 airfoil, a common engineering problem for assessing MF models [62, 63]. The aim is to predict the lift and drag coefficients ($C_{l}$ and $C_{d}$) considering two independent variables (freestream Mach number MA and angle of attack AoA). The dataset was obtained from [62], in which 5 HF points (simulated by Reynolds-averaged Navier Stokes equation) and 40 LF points (simulated by Euler equation) are provided for training, and 20 HF points are used for testing. Table 5 presents the results, and Figure 9 offers a comparison between all selected data-scarce MF methods.
Table 5: Results of different data-scarce methods for the NACA0012 airfoil MF problem for the two QoIs.
QoI Methods NRMSE R2 Score TLL Training time ($s$)
HF LF
Hierarchical Kriging 0.0058 0.9999 0.1636 0.0686 1.6828
Co-Kriging 0.0065 0.9999 4.2375 0.2138 1.4089
$C_{l}$ Scaled Kriging 0.0085 0.9999 3.3848 0.1270 2.1036
KRR-LR-GPR 0.0057 0.9999 4.3414 0.0258 1.0354
Hierarchical Kriging 0.1038 0.9107 5.9189 0.1259 1.2885
Co-Kriging 0.1089 0.9016 4.7347 0.1214 1.3689
$C_{d}$ Scaled Kriging 0.0987 0.9192 5.2907 0.0799 1.4836
KRR-LR-GPR 0.0923 0.9293 5.5690 0.0674 0.0745
Figure 9: Comparison results of the NACA0012 airfoil. The error bars of each point in the plot indicate predicted uncertainty
Table 5 shows that the performance of the proposed KRR-LR-GPR is better than other methods, especially when predicting $C_{d}$ but also for $C_{l}$ (observe the NRMSE values). The TLL value when predicting $C_{d}$ is the only metric that is marginally worse than Hierarchical Kriging. We note that KRR-LR-GPR is the fastest method to train, although this is not beneficial in this example because the HF and LF data acquisition time is much larger than the time needed to train any of the MF algorithms. In fact, this example is data-scarce even in the LF (only 40 points), and so the better scalability of KRR for LF regression is not needed in this case (this only becomes relevant when the LF dataset has over a few thousand points).
@@ 4.2.1 Illustrative example
We start with a one-dimensional illustrative example (data-scarce and low-dimensional) with 11 HF samples and 201 LF samples distributed uniformly. Aiming at explaining the limitations of the different MF models with DNNs and BNNs, we select three different LF datasets, labeled as “LF data 1”, “LF data 2” and “LF data 3”, obtained by establishing a different correlation with the HF data according to Equation 48b, (48c) and (48d), respectively. Model performance and accuracy metrics are shown in Figure 10 and Section 4.2.1.
@@
[h] Results comparison of DNN-LR-BNN on the 1D noisy illustrative example with 5 independent runs LF Dataset Methods NRMSE R2 Score TLL BNN 1.0470 (±$\pm 0.028$) -0.6888 (±$\pm 0.093$) -5.7825 (±$\pm 0.578$) LF Data 1 ($r$ = -0.1019) DNN-BNN 0.4865 (±$\pm 0.163$) 0.6029 (±$\pm 0.247$) 0.2808 (±$\pm 0.308$) DNN-LR1-BNN 1.0501 (±$\pm 0.013$) -0.6983 (±$\pm 0.043$) -6.4183 (±$\pm 0.662$) DNN-LR2-BNN 0.3161 (±$\pm 0.005$) 0.8461 (±$\pm 0.005$) 0.2329(±$\pm 0.090$) LF Data 2 ($r$ = 0.9999) DNN-BNN 0.3743 (±$\pm 0.092$) 0.7738 (±$\pm 0.102$) 0.5946 (±$\pm 0.136$) DNN-LR1-BNN 0.1051 (±$\pm 0.005$) 0.9830 (±$\pm 0.001$) 1.2296 (±$\pm 0.018$) LF Data 3 ($r$ = -0.0018) DNN-BNN 1.1012 (±$\pm 0.115$) -0.8834 (±$\pm 0.403$) -5.5084 (±$\pm 1.899$) DNN-LR1-BNN 1.0643 (±$\pm 0.012$) -0.7445 (±$\pm 0.038$) -6.1546 (±$\pm 0.711$)
    The subscripts 1 and 2 indicate that the LR model used for transfer-learning is order 1 (linear polynomial basis functions) and order 2 (quadratic polynomial basis functions), respectively.
The datasets LF data 1 and LF data 3 were deliberately created with a low Pearson correlation coefficient concerning the HF dataset. Figure 10 and Section 4.2.1 show that for such cases the LR transfer-learning model used in DNN-LR-BNN automatically finds low values for the $\rho_{1}$ coefficient after training the LF model on the HF data (independently of the BNN training on the residual). We believe that this is a strength of the proposed strategy because it automatically adjusts the influence of the LF model and informs the analyst about the correlation between LF and HF data. We note that the order of the LR transfer-learning model is a hyperparameter, but the coefficients $\boldsymbol{\rho}$ are found by training via minimizing Equation 6. Simple hyperparameter search strategies can quickly find the optimal order of the LR transfer-learning model for each problem (it is only one hyperparameter). For example, a quadratic LR transfer-learning model is better than a linear one for “LF data 1” (this is labeled as DNN-LR2-BNN to distinguish from the LR model with linear basis functions labeled as DNN-LR1-BNN). Evidently, if LF and HF data tend to be linearly correlated (as in the LF data 2), the DNN-LR-BNN model performs better.
@@
Figure 10: One-dimensional illustrative example comparing the proposed DNN-LR-BNN method (right column) with a BNN trained exclusively with HF data, and with a simple DNN-BNN (left column). Three LF datasets are employed whose Pearson correlation coefficients are $-0.1019$, $0.9999$, and $-0.0018$ correspondingly.
Interestingly, even when the correlation between LF data and HF data is highly nonlinear, we have not found the performance of the DNN-BNN to be better, despite this model using a BNN to perform transfer-learning instead of a simple LR model. The next sections demonstrate this for different datasets, including a material modeling example that illustrates a more realistic scenario encountered in engineering practice. This might explain why early MF literature considered first-order linear relationships between LF and HF models that were governed by a single hyperparameter ($\rho_{1}$). In essence, we find that performing model selection and establishing a simple (but not simpler) transfer-learning model between the DNN and BNN is an important consideration.
@@
We also want to bring to the reader’s attention that neural network training is highly dependent on architectural and hyperparametric choices. Section F.3 details the dependence on hyperparameters and reinforces that the DNN-LR-BNN model is simpler to train and is less sensitive to hyperparameter choices than the more flexible DNN-BNN model. We observed in all cases that the prediction of uncertainty depends on the width of the BNN, i.e. on the number of neurons in the hidden layers. A wide network leads to similar uncertainty predictions as the ones obtained by GPR, while the prediction of the expected value (the mean) converges quickly even for narrow BNNs.
4.2.2 Higher dimensional numerical examples
@@
Table 6 shows that the single-fidelity BNN leads to poor predictions for all examples according to all accuracy metrics. Also for all cases, the DNN-LR-BNN outperforms the DNN-BNN model. The coefficients $\hat{\bm{\rho}}$ learned for the DNN-LR-BNN model are reasonable when compared to the ground-truth functions shown in Section C.2.
4.2.3 Engineering application: material structure-property prediction dataset
This section focuses on finding structure-property relationships in materials, as described in [64]. The dataset considered here was introduced by Olivier et al. [65], where effective material properties of two-phase material microstructures are determined by computer simulations of material samples, called SVEs. Each SVE results from randomizing the material microstructure, then deforming the SVE according to periodic boundary conditions, and susbequently calculate the average (or homogenized) response of the SVE. The authors of that study only considered outputs from a single-fidelity, obtained by direct numerical analysis (DNS) via finite element analysis (FEA), and showed that a BNN can predict both the mean and standard deviation for chosen quantities of interest (QoIs). We extend this problem to the MF scenario by generating LF data using an open-source implementation [66] of a significantly faster simulation method called self-consistent clustering analysis (SCA) [7, 67]. Figure 11 summarizes the material structure-property regression problem, and additional information can be found in the original publication [65]. We highlight that a reader lacking domain knowledge of the Mechanics of Materials and multi-scale simulations may choose to ignore the physical meaning of the input and output variables of this dataset. In that case, the reader only needs to know that there are four input variables with the following bounds:
$V$ $\in[0.05,0.40],$
$E_{\text{fiber}}$ $\in[200000,600000],$
$a_{\text{matrix}}$ $\in[300,500],$
$b_{\text{matrix}}$ $\in[0.2,0.55],$
that lead to noisy measurements captured by three output variables: $E_{\text{hom}}$, $a_{\text{hom}}$ and $b_{\text{hom}}$, where the subscript “hom” indicates that these properties correspond to the homogenized (average) response of the two-phase material. The LF output measurements are obtained by a fast but less accurate method, while the HF measurements of the same three output variables are obtained by a slower but more accurate method. Note that both LF and HF measurements are noisy, as the same input point leads to different output values each time it is evaluated (due to material microstructure randomness). Figure 12 illustrates the variation of one of the outputs with respect to one of the inputs, while keeping the remaining inputs fixed. In that figure, red crosses correspond to HF data (DNS) and blue crosses to LF data (SCA).
Figure 11: Material structure-property prediction problem illustration. This problem involves four input descriptors: volume fraction $V$ of the particle material, Young’s modulus of the linear elastic fibers $E_{\texttt{fiber}}$, matrix hardening law coefficients $\sigma=400+a_{\text{matrix}}\varepsilon_{p}^{b_{\text{matrix}}}$. The simulations of the material response lead to three outputs: homogenized Young’s modulus $E_{\text{hom}}$, and homogenized hardening law with coefficients $a_{\text{hom}}$ and $b_{\text{hom}}$. We utilize two fidelity levels for data generation: HF data is generated through a FEA method via commercial software (ABAQUS) based on [68] and the LF data is obtained via the SCA method from the CRATE open source package [66]. Figures (a), (b), and (c) show three examples of how one output property changes concerning a particular input variable, making it clear that there is aleatoric uncertainty (noise). Section F.2 contains additional information about the data generation process for this problem.
We generated 500 HF samples with the FEA method and split them into 400 for training and 100 for testing. In addition to this, we also generate 3200 LF samples using the SCA method (using 3 material clusters – the hyperparameter of SCA that controls discretization). We then test the proposed DNN-LR-BNN strategy on this problem where we use 300 HF samples and 3200 LF samples to train the MF models, i.e. DNN-BNN and DNN-LR-BNN. We also train a single-fidelity BNN model using 400 HF samples, as this corresponds to the number of samples that can be generated with the same computational resources and time needed to create the MF dataset (in other words, 100 HF simulations generated with FEA require approximately the same time as 3200 LF simulations with the SCA method). Table 7 compares the performance of the different MF models, including the single-fidelity baseline of BNN. Figure 12 presents the output predictions compared to the measured output values in the test set. However, recall that the outputs are noisy and that their ground-truth mean values are not known, nor their corresponding aleatoric distribution (noise distribution). Therefore, the figure only shows the mean prediction (y-axis value), and corresponding epistemic uncertainty estimation (vertical bar) for each test sample (x-axis value).
Table 7: Results comparison of DNN-LR-BNN on the material structure-property linkages
QoIs Methods NRMSE R2 Score TLL $\hat{\rho}_{0}$ $\hat{\rho}_{1}$ $\hat{\rho}_{2}$
$a_{\text{eff}}$ ($r$ = 0.9294) BNN 0.0902 0.8165 -5.3577 - -
DNN-BNN 0.1254 0.6450 -5.2192 - -
DNN-LR1-BNN 0.0700 0.8894 -5.3748 10.00 0.92 -
DNN-LR2-BNN 0.0442 0.9558 -4.9248 10.00 1.24 -3.55 ×$\times 10^{-4}$
$b_{\text{eff}}$ ($r$ = 0.7681) BNN 0.0268 0.9489 3.2622 - -
DNN-BNN 0.0345 0.9155 3.1881 - -
DNN-LR1-BNN 0.0178 0.9774 3.3436 0.10 0.88 -
DNN-LR2-BNN 0.0134 0.9872 3.5999 0.05 1.11 0.32
$E_{\text{eff}}$ ($r$ = 0.9845) BNN 0.0025 0.9556 -10.8660 - -
DNN-BNN 0.0229 0.9649 -10.2588 - -
DNN-LR1-BNN 0.0146 0.9857 -9.2766 -7.97 1.01 -
DNN-LR2-BNN 0.0121 0.9901 -9.2447 -1.27 0.99 -7.75 ×$\times 10^{-8}$
Figure 12: Predicted outputs with uncertainties for material structure-property linkages problem. The size of the points represents the predicted uncertainty
Table 7 compares different MF models and also includes the single-fidelity BNN prediction. This table clarifies that the proposed DNN-LR-BNN model with a quadratic linear regression transfer learning model (DNN-LR2-BNN) has the best performance for this material structure-property problem. These results demonstrate the effectiveness of the DNN-LR-BNN method in real-world scenarios where the dataset contains multiple outputs. Figure 12 also highlights that the single-fidelity BNN has the largest predicted uncertainty for all test points. Conversely, the DNN-LR-BNN has a more reasonable uncertainty prediction, achieving the best TLL values among all methods. We can see that in cases where the outputs of the two fidelity levels have a high Pearson correlation coefficient ($r\rightarrow 1$), then the performance is similar for DNN-LR1-BNN and DNN-LR2-BNN because the training procedure identifies that the quadratic coefficient of the transfer-learning model in DNN-LR2-BNN tends to zero ($\hat{\rho}_{2}\rightarrow 0$). Instead, for the $b_{\text{hom}}$ output where the two fidelity levels are not as linearly correlated, the quadratic transfer learning model improves the prediction, as it facilitates training the BNN on the HF data (note that in this case we have $\hat{\rho}_{2}\neq 0$). We experienced difficulties in avoiding overfitting when training the DNN-BNN method for some of the outputs in this problem.
@@ 5 Discussion and limitations
Expressive machine learning algorithms typically have limited extrapolation ability. Their large number of parameters is used to explain the training data without necessarily discovering the underlying functional form that explains the data. DNNs often make unreliable predictions for points away from the support domain; conversely, Bayesian models such as GPRs and BNNs eventually reverse to the prior [69]. A similar issue also occurs in the DNN-BNN architecture where the HF prediction is distorted in the unsampled region as reported in Figure 13. However, KRR-LR-GPR and DNN-LR-BNN do not exhibit this behavior because they remain smooth and close to the HF truth. The proposed transfer-learning model in Equation 4 narrows the gap between $\mathbf{m}(\mathbf{X}^{h})^{T}\bm{\rho}$ and $\mathbf{y}^{h}$, while the BNN captures the residual utilizing a prior with zero mean that reverses to another constant after training (see purple line in Figure 13). Therefore, the HF prediction is influenced by the parsimonious (simple) transfer-learning model, improving its extrapolation ability.
Figure 13: Illustration of what occurs to the DNN-BNN, KRR-LR-GPR and DNN-LR-BNN model predictions when there is a domain region without HF samples ($x>1.0$), but considering LF samples throughout the entire domain ($20<x<$ 2).
However, when we compare the DNN-LR-BNN model prediction with the KRR-LR-GPR one, we see that although the mean is similar, the uncertainty prediction is not. The predicted uncertainty of DNN-LR-BNN is smaller in the extrapolation region – the model is overconfident. We identify two reasons for this observation. First, BNN inference is approximate and does not yield the exact posterior predictive distribution. This bottleneck can be resolved for a specific problem by fine-tuning the BNN hyperparameters using test data or cross-validation, as well as by increasing the number of neurons in the hidden layers. Second, there is an inherent bottleneck in evaluating uncertainty via Equation 14 when a BNN is employed for $f^{h}(\mathbf{x})$ inference: the covariance matrix $\mathbf{K(X,X)}$ is difficult to compute and so the uncertainty present in the LF model is difficult to transport to the HF prediction.